A cylinder with a base diameter of x units has a volume of πx3 cubic units.
Which statements about the cylinder are true? Check all that apply.
The radius of the cylinder is 1/2x units.
The radius of the cylinder is 2x units.
The area of the cylinder’s base is 1/4πx2 square units.
The area of the cylinder’s base is 1/2πx2 square units.
The height of the cylinder is 2x units.
The height of the cylinder is 4x units.

Answer by Boreal(15235)

Volume of a cylinder is pi*r^2h; r=(1/2)x, since the diameter is x units.
If height is 2x, then V=pi*(1/4)x^2*2x=pi(1/2)x^3, which is not pi*x^3
If height is 4x, then V=pi(1/4)x^2*4x=pi*x^3, which is true.
---------------
now exclude the others
radius is (1/2)x, not 1/2x or 2x. UNLESS, one means (1/2)x, which is true, but 1/(2x) is not true. The fraction needs to be clearer.
--------------
Area of the base is pi*r^2=pi (1/4)x^2; pi is in the numerator, so those two are excluded.
-------------